Simpliflying Square Roots
How to simply these:
1. 5 * 30
2. 64/48
3. 75
4. 35/7

Answer by stanbon(75887)

Simpliflying Square Roots
How to simply these:
1. sqrt(5*30 5sqrt(6)
2. sqrt(64/48) = sqrt(64)/sqrt(48)
= 8/4sqrt(3)
= 2/sqrt(3)
= 2sqrt(3)/3
= (2/3)sqrt(3)
----------------
3. sqrt(75)= 5sqrt(3)
4. sqrt(35/7) = sqrt(5)
=========================
Cheers,
Stan H.